OpenStudy (anonymous):

A fair die is tossed 4 times. Find the probability of a. at least 2 5s b. at most 3 even numbers c. no more than 1 odd number d. at least 1 prime number e. at most 1 number greater than 4

OpenStudy (yuki):

to find a), you need to know that P(at least 2 5's) = P(exactly 2 5s) + P(exactly 3 5s) +P(exactly 4 5s)+P(all 5s)

OpenStudy (yuki):

so you can find each of those probabilities and add them. that would be the straight forward way.

OpenStudy (yuki):

the easier way is to consider the complement of P(at least 5) which is P(at most 1)

OpenStudy (yuki):

that way all you have to do is calculate P(at most 1 5s) = P(no 5) + P(exactly 1 5)

OpenStudy (yuki):

Then you can use the fact that P(event A) = 1 - P( complement of event A)

OpenStudy (yuki):

does that help you at all ?

OpenStudy (yuki):

From the looks of it, all problems I see here becomes much easier if you use P(A) = 1-P(notA) let me know if you still need help

OpenStudy (anonymous):

how about the other parts?